(a) Q5, Chapter: Exchange Rates and the Foreign Exchange Market Calculate the dollar rates of ret urn on the following assets: i. A painting whose price rises from $200,000 to $250,000 in a year. ii. A bottle of a rare Burgundy, Domaine de la Romanée-Conti 2011 whose Drice rises from S255 to 875 ho on1
Answer #1

Annual rate of return = (Asset price in year 2 / Asset price in year 1) - 1

(i)

Rate of return = ($250,000 / $200,000) - 1 = 1.25 - 1 = 0.25 = 25%

(ii)

Rate of return = ($275 / $255) - 1 = 1.0784 - 1 = 0.0784 = 7.84%